# Part Segmentation of Human Meshes via Multi-View Human Parsing

### v2

@@ Abstract
Recent advances in point cloud deep learning have led to models that achieve high per-part labeling accuracy on large-scale point clouds, using only the raw geometry of unordered point sets. In parallel, the field of human parsing focuses on predicting body part and clothing/accessory labels from images. This work aims to bridge these two domains by enabling per-vertex semantic segmentation of large-scale human meshes. To achieve this, a pseudo-ground truth labeling pipeline is developed for the Thuman2.1 dataset: meshes are first aligned to a canonical pose, segmented from multiple viewpoints, and the resulting point-level labels are then backprojected onto the original mesh to produce per-point pseudo ground truth annotations. Subsequently, a novel, memory-efficient sampling strategy is introduced—windowed iterative farthest point sampling (FPS) with space-filling curve-based serialization—to effectively downsample the point clouds. This is followed by a purely geometric segmentation using PointTransformer, enabling semantic parsing of human meshes without relying on texture information. Experimental results confirm the effectiveness and accuracy of the proposed approach.
@@ 3 Method
An overview of the proposed pipeline for obtaining pseudo ground truth labels in Figure 3.1. Mesh re-orientation is detailed in Section 3.1, follwed by a description of the back-projection to 3D from 2D human parsing approaches in Section 3.2. The downsampling stage and overall segmentation approach in 3D is given in Section 3.3.
@@ 3.1 Mesh Orientation Alignment
The iterative alignment procedure operates by first rendering the mesh from a canonical viewpoint and extracting keypoint predictions using the trained pose estimator. The algorithm evaluates several geometric criteria to determine if further adjustment is necessary: (1) vertical orientation by comparing relative shoulder and hip positions, (2) depth orientation based on facial feature visibility (particularly nose confidence below 0.30.30.3 indicating a rear-facing pose), (3) lateral lean detection through shoulder confidence asymmetry exceeding 0.150.150.15, and (4) forward/backward lean assessment using eye confidence differentials.
When these conditions indicate misalignment, targeted rotations are applied: $\pi/2$ radians about the $z$-axis for horizontal corrections, $y$-axis sign inversion for vertical flips, $z$-axis sign inversion for depth corrections, and ±$\pm\pi/8$ or ±$\pm\pi/4$ radian rotations about the $y$-axis for lean adjustments based on relative keypoint confidences. The process iterates until either the pose metrics indicate proper alignment or a maximum iteration limit is reached, with pose improvement measured primarily through ear keypoint confidence enhancement above a threshold of 0.40.40.4, ensuring convergence toward the desired orientation where the person faces the camera in an upright stance.
@@ 3.2 Human Parsing and Backprojection
It is now assumed that the input mesh $M=\{V,E\}$ is properly aligned. The next step in the proposed approach is to compute human parsing segmentations for a set of views $v_{i}$ using pre-trained human parsing models. For each $v_{i}$, a parsing model is used which extracts a label image of size $I_{i}\in Z_{\geq 0}^{W\times H}$, with labels from a set $L=\{l_{1},l_{2},...,l_{L}\}$ with $l_{1}=0$ typically given to mean the background class. Views are sampled from elevations of {$0,-30\}$ , - 30 } and azimuth angles of $0\}$ , 90 , 180 , 270 }, both in degrees.
@@
[…]
@@ 3.3 Subsampling and Point Cloud Parsing Segmentation
In this section, it is assumed that there is a given dataset of point clouds with ground truth parsing labels $g_{j}$ for each vertices $n_{j}\in V$, for $N$ total points. For efficiency of preprocessing, training and inference, a novel sub-sampling approach will be employed, consisting first of Morton sorting (also referred to as serialization), followed by a windowed farthest point sampling algorithm. For each point cloud, the values are initially normalized to $[0,1]^{3}$ using coordinate-wise min-max normalization, yielding
@@
(removing left-trailing zeros). Points are sorted according to their Morton code and organized in $W$ windows of adjacent points, using padding of points with values at $[0,0,0]^{T}$ if necessary. Using the Morton approach, each window is locally coherent according to the properties of elementary space-filling curves, and by reducing the iterative farthest point sampling operation [17], for $k$ samples to $k/W$ samples for each of $k$ windows, run time is reduced from $O(kN)$ to $O(kN/W)$. Window size is chosen using a maximum number of points per window of 500050005000, where the total number of windows is
@@
The PointTransformer semantic segmentation model is organized into 111111 blocks of layers. Using K-NN as a feature aggregator, where for a point $x_{i}$, its features are computed as
@@ 6 Conclusion
Future research into this topic should largely focus on two main constituents. Firstly, the label space used for a general parsing approach should be as wide as possible, incorporating many clothing types and accessories that are not featured in the CIHP and Sapiens labels. This lack of label diversity can cause out of distribution issues in the closed-set model setting. Secondly, while the proposed approach can operate at inference time with under a second speed for an NVidia 4090 gpu (0.73$\approx$ 0.73≈ 0.73 seconds), this may still be too slow for certain applications. Morevoer, the number of views used as well as the use of the cpu-bound DBSCAN algorithm in the process of obtaining pseudo ground truth labels is slow and difficult to parallelize, hence more efficient methods should be pursued to speed up the overall pipeline.

### v3

@@ Abstract
Recent advances in point cloud deep learning have led to models that achieve high per-part labeling accuracy on large-scale point clouds, using only the raw geometry of unordered point sets. In parallel, the field of human parsing focuses on predicting body part and clothing/accessory labels from images. This work aims to bridge these two domains by enabling per-vertex semantic segmentation of large-scale human meshes. To achieve this, a pseudo-ground truth labeling pipeline is developed for the Thuman2.1 dataset: meshes are first aligned to a canonical pose, segmented from multiple viewpoints, and the resulting point-level labels are then backprojected onto the original mesh to produce per-point pseudo ground truth annotations. Subsequently, a novel, memory-efficient sampling strategy is introduced—windowed iterative farthest point sampling (FPS) with space-filling curve-based serialization—to effectively downsample the point clouds. This is followed by a purely geometric segmentation using PointTransformer, enabling semantic parsing of human meshes without relying on texture information. Experimental results confirm the effectiveness and accuracy of the proposed approach. Project code and pre-processed data is available at https://github.com/JamesMcCullochDickens/Human3DParsing/tree/master.
@@ 3 Method
An overview of the proposed pipeline for obtaining pseudo ground truth labels in Figure 3.1. Mesh re-orientation is detailed in Section 3.1, follwed by a description of the back-projection to 3D from 2D human parsing approaches in Section 3.2. The downsampling stage and overall segmentation approach in 3D is given in Section $3.$
@@ 3.1 Mesh Orientation Alignment
The iterative alignment procedure operates by first rendering the mesh from a canonical viewpoint and extracting keypoint predictions using the trained pose estimator. The algorithm evaluates several geometric criteria to determine if further adjustment is necessary: (1) vertical orientation by comparing relative shoulder and hip positions, (2) depth orientation based on facial feature visibility (particularly nose confidence below $0.3$ indicating a rear-facing pose), (3) lateral lean detection through shoulder confidence asymmetry exceeding $0.15$, and (4) forward/backward lean assessment using eye confidence differentials.
When these conditions indicate misalignment, targeted rotations are applied: $\pi/2$ radians about the $z$-axis for horizontal corrections, $y$-axis sign inversion for vertical flips, $z$-axis sign inversion for depth corrections, and $\pm\pi/8$ or $\pm\pi/4$ radian rotations about the $y$-axis for lean adjustments based on relative keypoint confidences. The process iterates until either the pose metrics indicate proper alignment or a maximum iteration limit is reached, with pose improvement measured primarily through ear keypoint confidence enhancement above a threshold of $0.4$, ensuring convergence toward the desired orientation where the person faces the camera in an upright stance.
@@ 3.2 Human Parsing and Backprojection
It is now assumed that the input mesh $M=\{V,E\}$ is properly aligned. The next step in the proposed approach is to compute human parsing segmentations for a set of views $v_{i}$ using pre-trained human parsing models. For each $v_{i}$, a parsing model is used which extracts a label image of size $I_{i}\in Z_{\geq 0}^{W\times H}$, with labels from a set $L=\{l_{1},l_{2},...,l_{L}\}$ with $l_{1}=0$ typically given to mean the background class. Views are sampled from elevations of $\{0,30,-30\}$ and azimuth angles of $\{0,90,180,270\}$, both in degrees.
@@
Following the voting procedure, the resulting vertex labels often contain noise and small disconnected components due to occlusions, rendering artifacts, and inconsistencies across views. To address this, a density-based spatial clustering approach using DBSCAN [34] is applied to refine the semantic assignments. For each semantic label class $l_{j}\in L$, vertices assigned to that class are extracted and subjected to DBSCAN clustering with parameters $\epsilon=0.03$ and $\text{min\_samples}=100$. The algorithm identifies the largest cluster as the primary component for each semantic class, while smaller clusters and noise points (vertices not belonging to any dense cluster) are relabeled using k-nearest neighbor (k-NN) propagation with $k=40$ neighbors from the remaining vertices. This denoising process effectively removes spurious label assignments and ensures spatial coherence in the final semantic segmentation, particularly important for preserving the anatomical structure of human body parts where disconnected components are typically undesirable. Further, a set of handcrafted rules are used per set of labels according to commonly observed artifacts. An example of the backprojection is shown in Figure $3.$
@@ 3.3 Subsampling and Point Cloud Parsing Segmentation
In this section, it is assumed that there is a given dataset of point clouds with ground truth parsing labels $g_{j}$ for each vertices $n_{j}\in V$, for NN total points. For efficiency of preprocessing, training and inference, a novel sub-sampling approach will be employed, consisting first of Morton sorting (also referred to as serialization), followed by a windowed farthest point sampling algorithm. For each point cloud, the values are initially normalized to $0,1]^{3}$ using coordinate-wise min-max normalization, yielding
@@
(removing left-trailing zeros). Points are sorted according to their Morton code and organized in $W$ windows of adjacent points, using padding of points with values at $0,0,0]^{T}$ if necessary. Using the Morton approach, each window is locally coherent according to the properties of elementary space-filling curves, and by reducing the iterative farthest point sampling operation [17], for $k$ samples to $k/W$ samples for each of $k$ windows, run time is reduced from $O(kN)$ to $O(kN/W)$. Window size is chosen using a maximum number of points per window of $5000$, where the total number of windows is
@@
The PointTransformer semantic segmentation model is organized into $11$ blocks of layers. Using K-NN as a feature aggregator, where for a point $x_{i}$, its features are computed as
@@ 6 Conclusion
Future research into this topic should largely focus on two main constituents. Firstly, the label space used for a general parsing approach should be as wide as possible, incorporating many clothing types and accessories that are not featured in the CIHP and Sapiens labels. This lack of label diversity can cause out of distribution issues in the closed-set model setting. Secondly, while the proposed approach can operate at inference time with under a second speed for an NVidia 4090 gpu ($\approx 0.73$ seconds), this may still be too slow for certain applications. Morevoer, the number of views used as well as the use of the cpu-bound DBSCAN algorithm in the process of obtaining pseudo ground truth labels is slow and difficult to parallelize, hence more efficient methods should be pursued to speed up the overall pipeline.